OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos^2(x) + 2cos(x) + 1 = 0

OpenStudy (anonymous):

I get cos(x)=-1 so the answers should be pi and 0 but the answer choices are: x = 2π x = π x = pi divided by four., seven pi divided by four. x = pi divided by two., three pi divided by two.

OpenStudy (jdoe0001):

well, you're correct, \(\bf cos^{-1}(cos(x)) = cos^{-1}(-1)\)

OpenStudy (anonymous):

\[\left( \cos x+1 \right)^{2}=0,\cos x=-1=\cos \pi \] \[x=\pi \]

OpenStudy (anonymous):

Since cos(x)=-1 the solution should have an x coordinate of -1, so shouldn't the solutions be pi, and 2pi

OpenStudy (anonymous):

no cos is negative in second and third quadrants only

OpenStudy (anonymous):

got it, thanks

OpenStudy (anonymous):

yw